OpenStudy (toxicsugar22):

a piece of copper has a mass of 150 gram? How many atoms does the sample contain.

OpenStudy (anonymous):

1

OpenStudy (anonymous):

\[150\text{ g}\cdot\frac{1\text{ mol}}{63.546\text{ g}}\cdot\frac{6.023\times10^{23}\text{ atoms}}{1\text{ mol}}=\cdots\]